Write the slope-intercept form of the equation of the line parallel to y = 5x + 2, which passes through the point (4, 9).
Answer by Cromlix(4381)

Lines that are parallel to one another
have the same gradient m1 = m2
Using the form y - b = m(x - a) and m = 5
y - 9 = 5(x - 4)
y - 9 = 5x - 20
y = 5x - 20 + 9
y = 5x - 11
Hope this helps.
:-)